# An exact analytical solution for the weakly magnetized flow around an axially symmetric paraboloid, with application to magnetosphere models

### v4

@@ Abstract
Rotationally symmetric bodies with longitudinal cross sections of parabolic shape are frequently used to model astrophysical objects, such as magnetospheres and other blunt objects, immersed in interplanetary or interstellar gas or plasma flows. We discuss a simple formula for the potential flow of an incompressible fluid around an elliptic paraboloid whose axis of symmetry coincides with the direction of incoming flow. Prescribing this flow, we derive an exact analytical solution to the induction equation of ideal magnetohydrodynamics for the case of an initially homogeneous magnetic field of arbitrary orientation being passively advected in this flow. Our solution procedure employs Euler potentials and Cauchy’s integral formalism based on the flow’s stream function and isochrones. Furthermore, we use a particular renormalization procedure that allows us to generate more general analytical expressions modeling the deformations experienced by arbitrary scalar or vector-valued fields embedded in the flow as they are advected first towards and then past the parabolic obstacle. Finally, both the velocity field and the magnetic field embedded therein are generalized from incompressible to mildly compressible flow, where the associated density distribution is found from Bernoulli’s principle.
@@ 1 Introduction
The study of fluid flow around solid obstacles is a classical topic of hydrodynamics, with countless applications in both engineering and physics, and especially astrophysics. Therein, the case of the fluid being magnetized is of particular interest for several classes of objects, ranging from the heliosphere (e.g. Röken et al., 2015; Kleimann et al., 2022), planets (e.g. Petrinec & Russell, 1997; Kotova et al., 2020) and their magnetospheres (e.g. Spreiter & Stahara, 1995; Richardson, 2002), comets and planetary satellites (e.g. Combi et al., 2002, and references therein), up to clusters of galaxies (e.g. Dursi & Pfrommer, 2008).
@@
This implies in particular that the flow field is fully determined by the shape of the obstacle and the flow conditions prevailing at the upstream boundary. In the case of planar symmetry, in which the shapes and flows are invariant along a Cartesian direction, classical examples of obstacles for which such potential flows have been explicitly derived include infinite planes, wedges, and the circular cylinder. By introducing a complex velocity potential, it is possible to generate additional solutions from the existing ones through the application of conformal mappings such as the Joukowski transform (e.g. Batchelor, 2000, pp. 435).
This paper, however, is concerned with settings where both the solid object and the flow around it feature an axial symmetry with respect to the inflow direction, rather than a translational one. For these, the method of complex potentials and conformal mappings is not available, and thus the set of known examples is correspondingly smaller.
@@
The set of known nontrivial solutions for magnetized generalizations of such flows is smaller still, having until now essentially been restricted to the sphere (e.g. Dursi & Pfrommer, 2008; Isenberg et al., 2015, and references therein) and the Rankine-type heliosphere (Röken et al., 2015; Isenberg et al., 2015). In the former case of a sphere, the exact magnetic field solution is known only up to an integral which so far could not be expressed in terms of analytical functions, whereas the latter Rankine-type case involves special functions, namely incomplete elliptic integrals. In contrast to these, the MHD solutions to be presented here are predominately of rather compact form and expressible through basic functions without recourse to unevaluated integrals, making them a viable alternative to numerical computations at least in simple applications.
Moreover, besides the obvious puristic and educational motivations, this geometrical setting is also of significant astrophysical interest. Starting with the work of Kobel & Flückiger (1994) and its subsequent refinements (e.g. Romashets & Vandas, 2019, and references therein), the upwind sides of planetary magnetopauses continue to be frequently modeled as rotationally symmetric paraboloids. Although the present approach cannot include discontinuous surfaces like bow shocks because it computes its velocity components from a smooth potential, it is certainly relevant at least in situations where a bow shock can either be neglected or is absent altogether.
@@
The paper is structured as follows. After this introduction, the analytical expressions and properties of the potential flow field around an axially-symmetric paraboloid are derived and discussed in Section 2. Thereafter, Section 3 presents the analytical derivation and discussion of a magnetic field $\vec{B}$ being passively advected in this flow by solving the stationary induction equation of ideal MHD
@@ 2 The inviscid flow field
In order to motivate the course of the present analysis, we start by considering the example of inviscid flow around a sphere, whose well-known flow potential and stream function (e.g. Batchelor, 2000, pp. 452) in terms of cylindrical coordinates $(\rho,\varphi,z)\in\mathbb{R}_{\geq 0}\times[0,2\pi)\times\mathbb{R}$ read, respectively,
@@
Since streamlines are given by lines of constant $\Psi_{0}$, we may interpret Equation (6) as a special case of the simpler stream function $\Psi_{\mathrm{bg}}:=-\rho^{2}/2$ of the undisturbed background flow $\vec{u}_{\mathrm{bg}}=-\vec{e}_{z}$ being modulated by a factor $1-1/f(\rho,z)$ which characterizes the shape of the obstacle in such a way that $f(\rho,z)=1$ represents its contour, in this case the sphere with $r=1$. The particular streamline $\Psi_{0}=0$ will then include not only the entire $z$-axis but split at two stagnation points (or one in cases where the obstacle is of infinite extent in the downstream direction), from which the emanating streamlines precisely trace the obstacle’s surface.
@@
of a conic section with its focal point at the origin, where $z=r\cos\vartheta$ and $\vartheta\in(0,\pi)$, covering ellipses ($10<\varepsilon<$ 1) and hyperbolas ($\varepsilon>1$) of eccentricity $\varepsilon$, as well as the special cases of circle ($\varepsilon=0$) and parabola ($\varepsilon=1$). But while the resulting flows around such ellipsoidal or hyperboloidal bodies of revolution are indeed solenoidal by construction (and might therefore be of actual interest for specialized applications), direct calculations show that no $g_{\varepsilon}(\cdot)$ that additionally satisfy Equation (1) exist for such bodies (see Appendix A). On the other hand, similar considerations can be used to generalize the sphere’s flow potential (5) to nonzero eccentricities, resulting in flow fields satisfying Equation (1) but not (2). We note that the correct components for flow around ellipsoids satisfying both constraints involve the solution of the Laplace equation $\nabla^{2}\Phi_{\varepsilon}=0$ in ellipsoidal coordinates (e.g. Milne-Thomson, 1968, pp. 475), and the rather complicated form of the associated expressions renders the general case $10<\varepsilon<$ 1 less attractive in the present context.
In the following, we will therefore focus on the parabolic case $\varepsilon=1$ with $g_{1}(r+z)=r+z$ being the identity mapping. This leads to stream function
@@
and the lower index ’1’, which was previously used to hint at $\varepsilon=1$, will be suppressed from now on. Equation (10) reproduces the stream function formula
@@
from Milne-Thomson (1968, p. 479), expressed in elliptical coordinates $(\xi,\eta)$ satisfying $(\rho,z)=\left(2\,c\,\xi\,\eta,[\xi^{2}-\eta^{2}]\,c\right)$ when choosing $c=U=1$ and $\xi_{0}=1/\sqrt{2}$, up to an overall minus sign (i.e. $\psi=-\Psi$) that originates from an alternative definition of the stream function used therein. Applying the relation
@@
which, by way of direct calculation, can indeed be shown to be irrotational (see again Appendix A). Therefore, a flow potential for this field exists, and is easily found to be of the simple form
@@
The flow field (14)–(15) is of course markedly simpler than the corresponding viscous flow that has been investigated analytically by Miller (1969, 1971) and numerically by, e.g. Davis & Werle (1972) because it avoids the intricate modeling of boundary layers that is required when solving the Navier–Stokes equation for the geometry in question.
@@ 3.1 The longitudinal and transversal components
Next, we turn to the computation of the components of a magnetic field which is embedded into the flow (14)–(15) and passively advected in it. As often in such cases, we require the magnetic field at upstream infinity ($z\rightarrow+\infty$) to be of constant (though arbitrary) orientation and magnitude. The boundary field is best specified through its constant Cartesian components $(B_{x0},B_{y0},B_{z0})$, yielding the cylindrical representation
@@
of the magnetic field into a longitudinal (flow-parallel) component $\vec{B}_{\mathrm{lo}}$ and a transversal component $\vec{B}_{\mathrm{tr}}$, which at the upstream boundary (but generally not at other locations) is perpendicular to the flow. When introducing the azimuthal angle $\varphi_{0}$ of the boundary field, i.e. the unique angle satisfying
@@
using Euler potentials $\alpha$ and $\beta$ (e.g. Stern, 1966), which are scalar functions of position. This ansatz implies in particular that field lines may be identified as the lines at which surfaces of constant $\alpha$ intersect those of constant $\beta$. Furthermore, it guarantees that the divergence constraint $\nabla\cdot\vec{B}_{\mathrm{tr}}=0$ is always satisfied. Here, we choose the first Euler potential $\alpha$ as
@@ 3.2 Isochrones
The second Euler potential $\beta$ is chosen as the negative of the travel time $t$ that a fluid particle spent moving from upstream infinity along a trajectory $t\mapsto(\rho(t),z(t))$ to its current position, and lines (or, in three dimensions, surfaces) of constant $\beta$ will therefore be referred to as isochrones. The minus sign, which is absent from the formulas used by Isenberg et al. (2015), has to be added to either $\alpha$ or $\beta$ to ensure consistency with the desired boundary conditions (17)–(19), and we chose $\beta$ for reasons that will become clear in Section 4.2.
@@
The symbol $\Psi$ in the integrand indicates that integration has to be performed along a streamline, i.e. at constant $\Psi$, and $a$ is the value attained by the $\rho$ coordinate of the streamline in question at $z\rightarrow\infty$, that is, the value satisfying
@@
We point out that Equation (28) suffers from the conceptual problem that it will necessarily lead to infinite travel times for any $(\rho,z)$, consistent with an infinitely long distance being traversed at finite speed. Actual computations, such as that of the resulting magnetic field components via Equation (25), therefore require the integral to start at a finite height $z_{0}$ (and a corresponding axial distance $\rho_{0}$) to keep the travel time finite. Only afterwards may one perform the limit $(\rho_{0},z_{0})\rightarrow(a,\infty)$ along the streamline in question. (See Section 3.4 for a more sophisticated remedy.)
@@
$=\frac{1}{2}\ln\bigg{(}\frac{\sqrt{(r-z)\,(r+z-1)+(z_{0}-1/2)^{2}}}{r-z}-\bigg{.}$
@@
$\nabla\beta=$ $\lim_{z_{0}\rightarrow\infty}\nabla\hat{\beta}=\frac{\rho}{2\,(r+z-1)\,r}\,\vec{e}_{\rho}+$
@@
$\frac{(B_{\mathrm{tr}})_{\rho}}{B_{\mathrm{tr,0}}}=$ $\ \cos(\varphi-\varphi_{0})\,\sqrt{\frac{r+z}{r+z-1}}\times$
@@ 3.3 General field structure
implying that in the limit $z\rightarrow 1/2$ (i.e. at the stagnation point, located where the $z$-axis intersects the obstacle’s surface), the longitudinal component vanishes while the transversal component tends to infinity. This behavior is analogous to what is found for the Rankine heliosphere, where the corresponding term in round brackets reads $(1-1/z^{2})$ and likewise tends to infinity as $z\rightarrow 1$ (see Appendix C of Röken et al., 2015). Indeed, it can be verified that this divergent field strength is not restricted to the stagnation point but extends to the entire surface of the paraboloid, also in exact analogy to the heliosphere case. The reason for this can be traced to a mismatch between the conflicting assumptions of stationarity and idealness, which enforces a perpetual pile-up of magnetic flux that cannot be dissipated away. In a more realistic nonideal setting, the rate of incoming advected flux would be balanced by resistive dissipation, allowing for an equilibrium state with finite field strength. This, however, is beyond the scope of the present approach, which crucially relies on the frozen-in condition that is ensured by the ideal induction equation (4). A viable option to mitigate this problem for the heliopause, which would likely work also in the present case, has been presented by Florinski et al. (2024).
@@ 3.4 Isochrones revisited
For the applications to be discussed in Section 4, it would be convenient to have at our disposal a set of isochrones which, while leading to the same magnetic field, are finite-valued in the entire region of interest. We will now introduce a procedure to arrive at such an isochrone field.

### v5

@@ Abstract
Rotationally symmetric bodies with longitudinal cross sections of parabolic shape are frequently used to model astrophysical objects, such as magnetospheres and other blunt objects, immersed in interplanetary or interstellar gas or plasma flows. We discuss a simple formula for the potential flow of an incompressible fluid around an elliptic paraboloid whose axis of symmetry coincides with the direction of incoming flow. Prescribing this flow, we derive an exact analytical solution to the induction equation of ideal magnetohydrodynamics for the case of an initially homogeneous magnetic field of arbitrary orientation being passively advected in this flow. Our solution procedure employs Euler potentials and Cauchy’s integral formalism based on the flow’s stream function and isochrones. Furthermore, we use a particular renormalization procedure that allows us to generate more general analytical expressions modeling the deformations experienced by arbitrary scalar or vector-valued fields embedded in the flow as they are advected first toward and then past the parabolic obstacle. Finally, both the velocity field and the magnetic field embedded therein are generalized from incompressible to mildly compressible flow, where the associated density distribution is found from Bernoulli’s principle.
@@ 1 Introduction
The study of fluid flow around solid obstacles is a classical topic of hydrodynamics, with countless applications in both engineering and physics, and especially astrophysics. Therein, the case of the fluid being magnetized is of particular interest for several classes of objects, ranging from the heliosphere (e.g., Röken et al., 2015; Kleimann et al., 2022), planets (e.g., Petrinec & Russell, 1997; Kotova et al., 2020) and their magnetospheres (e.g., Spreiter & Stahara, 1995; Richardson, 2002), comets and planetary satellites (e.g., Combi et al., 2002, and references therein), up to clusters of galaxies (e.g., Dursi & Pfrommer, 2008).
@@
This implies in particular that the flow field is fully determined by the shape of the obstacle and the flow conditions prevailing at the upstream boundary. In the case of planar symmetry, in which the shapes and flows are invariant along a Cartesian direction, classical examples of obstacles for which such potential flows have been explicitly derived include infinite planes, wedges, and the circular cylinder. By introducing a complex velocity potential, it is possible to generate additional solutions from the existing ones through the application of conformal mappings such as the Joukowski transform (e.g., Batchelor, 2000, p. 435).
@@
This paper, however, is concerned with settings where both the solid object and the flow around it feature an axial symmetry with respect to the inflow direction, rather than a translational one. For these, the method of complex potentials and conformal mappings is not available, and thus the set of known examples is correspondingly smaller. The set of known nontrivial solutions for magnetized generalizations of such flows is smaller still, having until now essentially been restricted to the sphere (e.g., Dursi & Pfrommer, 2008; Isenberg et al., 2015, and references therein) and the Rankine-type heliosphere (Röken et al., 2015; Isenberg et al., 2015). In the former case of a sphere, the exact magnetic field solution is known only up to an integral which so far could not be expressed in terms of analytical functions, whereas the latter Rankine-type case involves special functions, namely, incomplete elliptic integrals. In contrast to these, the MHD solutions to be presented here are predominately of rather compact form and expressible through basic functions without recourse to unevaluated integrals, making them a viable alternative to numerical computations at least in simple applications.
Moreover, besides the obvious puristic and educational motivations, this geometrical setting is also of significant astrophysical interest. Starting with the work of Kobel & Flückiger (1994) and its subsequent refinements (e.g., Romashets & Vandas, 2019, and references therein), the upwind sides of planetary magnetopauses continue to be frequently modeled as rotationally symmetric paraboloids. More generally, the situation of a flow impacting on a solid, convex object is obviously a very generic one. It is thus also potentially applicable to the interaction of other objects like astrospheres and their galactic analogs with their respective magnetized environments. Particularly the latter has received relatively little attention since the works of Ikeuchi & Tomisaka (1981) but is becoming increasingly relevant in the light of more recent observations and associated modeling of cosmic rays accelerated at a Galactic termination shock (Merten et al., 2018), as well as works on galaxy-scale interactions such as ram-pressure stripping (e.g., Ramos-Martínez et al., 2018; Müller et al., 2021).
It is clear that the numerous simplifying assumptions inherent to analytical approaches of the present kind, such as the stationarity and homogeneity of the incident flow, the perfectly regular shape of the geometric surface, and the absence of turbulent small-scale motion, will typically render them less suitable for an accurate reproduction of real-world astrophysical scenarios, which are therefore better left to numerical investigations. Then again, it is exactly this simplifying approach that allows for the model to be applied to such a relatively wide range of different (astro)physical scenarios in the first place. And while the present approach cannot include discontinuous surfaces like bow shocks because it computes its velocity components from a smooth potential, it is certainly relevant at least in situations where a bow shock can either be neglected or is absent altogether.
@@
The paper is structured as follows. After this introduction, the analytical expressions and properties of the potential flow field around an axially symmetric paraboloid are derived and discussed in Section 2. Thereafter, Section 3 presents the analytical derivation and discussion of a magnetic field $\vec{B}$ being passively advected in this flow by solving the stationary induction equation of ideal MHD,
@@ 2 The inviscid flow field
In order to motivate the course of the present analysis, we start by considering the example of inviscid flow around a sphere, whose well-known flow potential and stream function (e.g., Batchelor, 2000, p. 452) in terms of cylindrical coordinates $(\rho,\varphi,z)\in\mathbb{R}_{\geq 0}\times[0,2\pi)\times\mathbb{R}$ read, respectively,
@@
Since streamlines are given by lines of constant $\Psi_{0}$, we may interpret Equation (6) as a special case of the simpler stream function $\Psi_{\mathrm{bg}}:=-\rho^{2}/2$ of the undisturbed background flow $\vec{u}_{\mathrm{bg}}=-\vec{e}_{z}$ being modulated by a factor $1-1/f(\rho,z)$ that characterizes the shape of the obstacle in such a way that $f(\rho,z)=1$ represents its contour, in this case the sphere with $r=1$. The particular streamline $\Psi_{0}=0$ will then include not only the entire $z$-axis but split at two stagnation points (or one in cases where the obstacle is of infinite extent in the downstream direction), from which the emanating streamlines precisely trace the obstacle’s surface.
@@
of a conic section with its focal point at the origin, where $z=r\cos\vartheta$ and $\vartheta\in(0,\pi)$, covering ellipses ($10<\varepsilon<$ 1) and hyperbolas ($\varepsilon>1$) of eccentricity $\varepsilon$, as well as the special cases of circle ($\varepsilon=0$) and parabola ($\varepsilon=1$). However, while the resulting flows around such ellipsoidal or hyperboloidal bodies of revolution are indeed solenoidal by construction (and might therefore be of actual interest for specialized applications), direct calculations show that no $g_{\varepsilon}(\cdot)$ that additionally satisfies Equation (1) exists for such bodies (see Appendix A). On the other hand, similar considerations can be used to generalize the sphere’s flow potential (5) to nonzero eccentricities, resulting in flow fields satisfying Equation (1) but not (2). We note that the correct components for flow around ellipsoids satisfying both constraints involve the solution of the Laplace equation $\nabla^{2}\Phi_{\varepsilon}=0$ in ellipsoidal coordinates (e.g., Milne-Thomson, 1968, p. 475), and the rather complicated form of the associated expressions renders the general case $10<\varepsilon<$ 1 less attractive in the present context.
In the following, we will therefore focus on the parabolic case $\varepsilon=1$ with $g_{1}(r+z)=r+z$ being the identity mapping. This leads to the stream function
@@
and the lower index “1,” which was previously used to hint at $\varepsilon=1$, will be suppressed from now on. Equation (10) reproduces the stream function formula
@@
from Milne-Thomson (1968, p. 479), expressed in elliptical coordinates $(\xi,\eta)$ satisfying $(\rho,z)=\left(2\,c\,\xi\,\eta,[\xi^{2}-\eta^{2}]\,c\right)$ when choosing $c=U=1$ and $\xi_{0}=1/\sqrt{2}$, up to an overall minus sign (i.e., $\psi=-\Psi$) that originates from an alternative definition of the stream function used therein. Applying the relation
@@
which, by way of direct calculation, can indeed be shown to be irrotational (see again Appendix A). Therefore, a flow potential for this field exists and is easily found to be of the simple form
@@
The flow field (14)–(15) is of course markedly simpler than the corresponding viscous flow that has been investigated analytically by Miller (1969, 1971) and numerically by, e.g., Davis & Werle (1972) because it avoids the intricate modeling of boundary layers that is required when solving the Navier–Stokes equation for the geometry in question.
@@ 3.1 The longitudinal and transversal components
Next, we turn to the computation of the components of a magnetic field that is embedded into the flow (14)–(15) and passively advected in it. As often in such cases, we require the magnetic field at upstream infinity ($z\rightarrow+\infty$) to be of constant (though arbitrary) orientation and magnitude. The boundary field is best specified through its constant Cartesian components $(B_{x0},B_{y0},B_{z0})$, yielding the cylindrical representation
@@
of the magnetic field into a longitudinal (flow-parallel) component $\vec{B}_{\mathrm{lo}}$ and a transversal component $\vec{B}_{\mathrm{tr}}$, which at the upstream boundary (but generally not at other locations) is perpendicular to the flow. When introducing the azimuthal angle $\varphi_{0}$ of the boundary field, i.e., the unique angle satisfying
@@
using Euler potentials $\alpha$ and $\beta$ (e.g., Stern, 1966), which are scalar functions of position. This ansatz implies in particular that field lines may be identified as the lines at which surfaces of constant $\alpha$ intersect those of constant $\beta$. Furthermore, it guarantees that the divergence constraint $\nabla\cdot\vec{B}_{\mathrm{tr}}=0$ is always satisfied. Here, we choose the first Euler potential $\alpha$ as
@@ 3.2 Isochrones
The second Euler potential $\beta$ is chosen as the negative of the travel time $t$ that a fluid particle spent moving from upstream infinity along a trajectory $t\mapsto\big{(}\rho(t),z(t)\big{)}$ to its current position, and lines (or, in three dimensions, surfaces) of constant $\beta$ will therefore be referred to as isochrones. The minus sign, which is absent from the formulas used by Isenberg et al. (2015), has to be added to either $\alpha$ or $\beta$ to ensure consistency with the desired boundary conditions (17)–(19), and we chose $\beta$ for reasons that will become clear in Section 4.2.
@@
The symbol $\Psi$ in the integrand indicates that integration has to be performed along a streamline, i.e., at constant $\Psi$, and $a$ is the value attained by the $\rho$ coordinate of the streamline in question at $z\rightarrow\infty$, that is, the value satisfying
@@
We point out that Equation (28) suffers from the conceptual problem that it will necessarily lead to infinite travel times for any $(\rho,z)$, consistent with an infinitely long distance being traversed at finite speed. Actual computations, such as that of the resulting magnetic field components via Equation (25), therefore require the integral to start at a finite height $z_{0}$ (and a corresponding axial distance $\rho_{0}$) to keep the travel time finite. Only afterward may one perform the limit $(\rho_{0},z_{0})\rightarrow(a,\infty)$ along the streamline in question. (See Section 3.4 for a more sophisticated remedy.)
@@
$=\frac{1}{2}\ln\bigg{(}\frac{\sqrt{(r-z)\,(r+z-1)+(z_{0}-1/2)^{2}}}{r-z}\bigg{.}$
@@
$\nabla\beta=$ $\lim_{z_{0}\rightarrow\infty}\nabla\hat{\beta}=\frac{\rho}{2\,(r+z-1)\,r}\,\vec{e}_{\rho}$
@@
$\frac{(B_{\mathrm{tr}})_{\rho}}{B_{\mathrm{tr,0}}}=$ $\ \cos(\varphi-\varphi_{0})\,\sqrt{\frac{r+z}{r+z-1}}$
@@ 3.3 General field structure
implying that in the limit $z\rightarrow 1/2$ (i.e., at the stagnation point, located where the $z$-axis intersects the obstacle’s surface), the longitudinal component vanishes while the transversal component tends to infinity. This behavior is analogous to what is found for the Rankine heliosphere, where the corresponding term in round brackets reads $(1-1/z^{2})$ and likewise tends to zero as $z\rightarrow 1$ (see Appendix C of Röken et al., 2015). Indeed, it can be verified that the resulting divergent field strength is not restricted to the stagnation point but extends to the entire surface of the paraboloid, also in exact analogy to the heliosphere case. The reason for this can be traced to a mismatch between the conflicting assumptions of stationarity and idealness, which enforces a perpetual pileup of magnetic flux that cannot be dissipated away. In a more realistic nonideal setting, the rate of incoming advected flux would be balanced by resistive dissipation, allowing for an equilibrium state with finite field strength. This, however, is beyond the scope of the present approach, which crucially relies on the frozen-in condition that is ensured by the ideal induction equation (4). A viable option to mitigate this problem for the heliopause, which would likely work also in the present case, has been presented by Florinski et al. (2024).
@@ 3.4 Isochrones revisited
For the applications to be discussed in Section 4, it would be convenient to have at our disposal a set of isochrones that, while leading to the same magnetic field, are finite-valued in the entire region of interest. We will now introduce a procedure to arrive at such an isochrone field.